OpenStudy (anonymous):

if one zero is polynomial f(x)=(kxk+4)XxX+13x+4k is reciprocal of the others,then k=?

sam (.sam.):

What's kxk and XxX

OpenStudy (raden):

f(x)= (k^2 + 4)x^2 + 13x + 4k, is it like this ?

OpenStudy (raden):

if so, x+1/x will satisfies when a = c with a coeff of x^2 and c is constant

OpenStudy (raden):

known, a = k^2 + 4, and c = 4k a = c k^2 + 4 = 4k k^2 - 4k + 4 = 0 solve for k ....

OpenStudy (raden):

@Manishakumari